This thread has been locked.
If you have a related question, please click the "Ask a related question" button in the top right corner. The newly created question will be automatically linked to this question.
Part Number: TPD12S016
We just had some HDMI compliance tests run on our unit which incorporates your TPB12S016. We failed DDC/CEC Line Capacitance and Voltage. The Capacitance is too high 650p on SDA and 950p on SCL.
While I do not have a LCR bridge that can be set per the test procedure (bias 2.5V, AC voltage 3.5V, Freq 100K) using my LCR bridge (10K, 1V, no bias) I measure a total Jig+unit capacitance of ~30p without the chip, and 120/190p with the chip (Jig capacitance ~12p).
We are connected as a sink, SDA/SCL go to port B. My interpretation of your data sheet shows input C at 15p. My measured connector/track capacitance is ~18p, so 18p + your 15p would be 33p and within the speced 50p max...
Any feedback would be helpful
We are glad that we were able to resolve this issue, and will now proceed to close this thread.
If you have further questions related to this thread, you may click "Ask a related question" below. The newly created question will be automatically linked to this question.
In reply to Chuck Branch:
Do you see anything improper with the way we have the part connected?
In reply to Ed Judziewicz:
I finally have the unit we sent in for testing.
When we measure the capacitance at 20KHz or below. We get ~30pF which makes sense. But above say 30K we see a glitch in the stimulus waveform. See attached. I think this is probably giving false readings on the test gear.
TI TPD12S016 100K stimulas 3.5Vp-p, 2.5V offset.docx
For the HDMI test we need to supply a 3.5Vp-p 100K stimulus signal.
Since I do not have the "official" setup... What we did to try and test this was to feed a 7Vp-p, 100K sine, centered on ground through a 1.5K resistor to the I2C pin under test. This creates a divider via the internal1.5K pull up to 5V and presents a 3.5Vp-p (ish) signal biased at 2.5V.
We did NOT see the switching feedback glitch at frequencies below 20KHz. Above 45KHz the signal appears and eventually merges into one pulse above100KHz.
There is mention of a "rise time accelerator" How does this impact the test?
At least two parts responded the same.
I think you are onto something with looking at the rist time accelerator. I will attempt to contact the design team to get a definitive answer, but your captures provide good evidence that the LRC measurement is being corrupted by the TPD12S016.
Can you run the same test where your input sine wave is dc shifted so that the input to the TPD12S016 is always above 800mV. This should prevent the acceleration circuit from turning on.
Shifting the sine wave did not have an impact.
Reducing amplitude idi not help
Any response from the design team?
All content and materials on this site are provided "as is". TI and its respective suppliers and providers of content make no representations about the suitability of these materials for any purpose and disclaim all warranties and conditions with regard to these materials, including but not limited to all implied warranties and conditions of merchantability, fitness for a particular purpose, title and non-infringement of any third party intellectual property right. No license, either express or implied, by estoppel or otherwise, is granted by TI. Use of the information on this site may require a license from a third party, or a license from TI.
TI is a global semiconductor design and manufacturing company. Innovate with 100,000+ analog ICs andembedded processors, along with software, tools and the industry’s largest sales/support staff.